### v1

# Common Inpainted Objects In-N-Out of Context

Abstract
We present Common Inpainted Objects In-N-Out of Context (COinCO), a novel dataset addressing the scarcity of out-of-context examples in existing vision datasets. By systematically replacing objects in COCO images through diffusion-based inpainting, we create 97,722 unique images featuring both contextually coherent and inconsistent scenes, enabling effective context learning. Each inpainted object is meticulously verified and categorized as in- or out-of-context through a multimodal large language model assessment. Our analysis reveals significant patterns in semantic priors that influence inpainting success across object categories. We demonstrate three key tasks enabled by COinCO: (1) training context classifiers that effectively determine whether existing objects belong in their context; (2) a novel Objects-from-Context prediction task that determines which new objects naturally belong in given scenes at both instance and clique levels, and (3) context-enhanced fake detection on state-of-the-art methods without fine-tuning. COinCO provides a controlled testbed with contextual variations, establishing a foundation for advancing context-aware visual understanding in computer vision and image forensics. Our code and data are at: https://github.com/YangTianze009/COinCO.
1 Introduction
Context is fundamental to visual understanding [1, 2, 3]. When humans view a scene, we instinctively assess the contextual coherence between objects and their environment. This context-based reasoning is essential for interpreting real-world scenes [4] and beyond. Take a look at Figure 2—can you spot the inpainted (i.e., fake) object in each image? Contextual understanding helps identify objects that violate scene expectations, even when pixel-level artifacts are imperceptible. By using context, humans naturally assess whether objects appear in plausible settings—a potted plant flying in the sky would immediately raise suspicion, while the same plant in a garden appears perfectly natural.
Learning context from data, however, is difficult. A significant challenge is that unusual scenes with out-of-context objects are rare in real life. Common computer vision datasets [5, 6, 7, 8] primarily contain objects in their natural settings, creating a scarcity of examples with contextual violations. This presents a fundamental obstacle: how can we train machine learning models that require large data to recognize contextual inconsistencies when such examples are inherently uncommon?
To address this challenge, we propose a dataset with both in-context and out-of-context objects through controlled image manipulation. By systematically replacing objects in existing real scenes, we can generate the necessary examples of contextual violations while preserving the overall structure of the scene. We specifically chose the Common Objects in Context (COCO) dataset [5] as our foundation because it contains a diverse set of everyday scene photographs with rich image- and object-level annotations, offering an ideal starting point for contextual manipulation.
Figure 1: Which object is fake? Only one object per image is inpainted. Out-of-context inpainted objects are easier to identify. Answers are revealed at the bottom of this page [^2] Figure 2 answers. 1st row: bird, laptop; 2nd row: hotdog, potted plant, potted plant; 3rd row: horse, horse..
Using Stable Diffusion’s inpainting model [9], we replace exactly one object per COCO image. This selective approach allows us to maintain the broader scene context while introducing precise, controlled variations in object-scene relationships. Our meticulous data creation pipeline and multi-step verification ensure high-quality inpainting results, and we apply a state-of-the-art Multimodal Large Language Model (MLLM) to classify each inpainted object as either contextually consistent or inconsistent with its scene.
Figure 2 shows the pipeline of our data creation and the downstream tasks. We name our dataset Common Inpainted Objects In-N-Out of Context (COinCO), reflecting our emphasis on context for inpainted objects. Our dataset contains 97,722 unique inpainted images, each accompanied by rich annotations.
COinCO is the first dataset that features inpainted objects with diverse context annotations, enabling advancements in tasks including context classification, object-from-context prediction, and fake detection (as our objects are inpainted by a generative model). COinCO serves as a challenging benchmark, encouraging new research for the computer vision community.
The main contributions of our work are:
1. We introduce a novel, large-scale dataset of partially manipulated images that enhances COCO, featuring strategically inpainted objects that are either contextually coherent or inconsistent with their scenes.
2. We employ COinCO’s context labels to train a context classifier. This task utilizes the visual features of an image, alongside the semantic features of detected objects to predict an object to be in- or out-of-context.
3. We introduce a novel task, object-from-context, which aims to predict instance- and clique-level categories for possible objects that fit the given context.
4. We demonstrate that context can be effectively integrated into fake detection pipelines, substantially improving state-of-the-art methods without fine-tuning.
2 Related Work
COCO and its extensions. The Common Objects in Context (COCO) dataset [5] is a widely used benchmark for various computer vision tasks, including object detection, instance segmentation, and captioning. Over the years, several extensions have been proposed. Some datasets primarily focus on expanding annotations without modifying image content. COCO-WholeBody refines human keypoint detection by adding facial, hand, and foot keypoints [10], while LVIS introduces a long-tailed distribution with over 1,000 categories for instance segmentation [7]. COCO-Stuff incorporates background (stuff) annotations for panoptic segmentation [6], and RefCOCO enables referring expression-based object localization [8]. Other works improve vision-language alignment, such as COCO-Caption for image captioning [11] and COCO-Text for scene text detection [12]. Additionally, CD-COCO, which applies image distortions to test robustness [13], and COCONut, which unifies multiple segmentation tasks [14], explore scene complexity. However, these datasets do not manipulate contextual relationships. Our dataset uniquely reconstructs scene context by replacing original objects with out-of-context alternatives via inpainting, creating a testbed for context modeling and fake localization that enables the learning of complex scene semantics.
Context reasoning. Context is essential for understanding object relationships and finding anomalies in complex scenes. Biederman et al. [4] identified five relational principles—support, interposition, probability, position, and size—that help identify contextual inconsistencies. Prior works apply these principles using object co-occurrence and support relationships for out-of-context detection [2]. Others have demonstrated that slight context changes in “object transplanting” experiments can cause significant errors in object detection. Acharya et al. [15] introduced a Graph Contextual Reasoning Network (GCRN) to model co-occurrence and relative position to detect out-of-context objects. A comprehensive review of context in vision is in [16]. Ours is the first work to feature a mixture of in- and out-of-context objects in large-scale data. Moreover, context is under-explored in fake detection. Multimodal Large Language Models (MLLMs)  [17, 18, 19, 20, 21], offer a human-aligned approach to nuanced context reasoning. We are the first to use context in fake detection with MLLMs.
Figure 2: Our COinCO pipeline. (a) For a given COCO image, an object is randomly replaced by with Stable Diffusion inpainting. (b) Inpainting success is verified using object detection and an MLLM. Successes are added to the dataset, while fail cases are regenerated and retested. (c) Inpainted images are classified as in-context or out-of-context using the MLLM. (d) Instance-level (object category) and clique-level (supercategory) information is predicted from the inpainted image using an MLP and recovered with Stable Diffusion. (e) Fake pixels are detected with pretrained fake detectors and enhanced within the bounding box region of a detected out-of-context object. (f) Semantic and visual embeddings are derived from the object list and inpainted image, respectively. These embeddings are used to train an MLP for classifying instances as in- or out-of-context.
Fake image generation and detection. Image manipulation has a long history from traditional object insertion techniques [22, 23] to GAN-based synthesis [24, 25, 26, 27, 28]. Recently, diffusion models set new standards [29, 30], with text-to-image models like Stable Diffusion [9] and DALL-E [31] ControlNet enhances diffusion models with controls [32]. Emerging techniques tailor diffusion methods for object manipulation, inpainting, and harmonization [33, 34], 3D [35], and relighting [36, 37].
For detecting manipulated images, earlier works classify at image-level [38, 39]. Advanced methods detect various manipulations with details: PSCC-Net [40] employs spatio-channel correlation across scales, CAT-Net [41] focuses on JPEG compression artifacts, and ManTra-Net [42] captures diverse manipulation traces. TruFor combines RGB and noise-sensitive fingerprints to produce anomaly and confidence maps [43]. For fake image datasets, e.g., GenImage [44], CIFAKE [45], DE-FAKE [30], they focus on fully-synthesized images. Diffusion-based inpainting datasets like COCOGlide [43] and TGIF [46] utilize COCO images but they have limited image quantity, object replacements constrained to the same category (thus lacking out-of-context objects), and no context labels. Our dataset contains significant out-of-context fake objects in real scenes with rich characteristics.
3 Building COinCO
We propose COinCO, a novel context-oriented inpainted objects dataset derived from the widely-used Common Objects in Context (COCO) Dataset [5]. Leveraging COCO’s image- and object-level annotations, COinCO enhances COCO by systematically replacing objects with diffusion-based inpainting, providing images with enriched contextual diversity. This rendition has 95,320 unique training images and 2,402 testing images from the COCO2017 training and validation sets, respectively. Each inpainted object is categorized as in- or out-of-context and is accompanied by information about the original object and its replacement.
For each COCO image, we (1) randomly pick an object from the image and perform inpainting using Stable Diffusion, (2) verify the new object was successfully inpainted with object detection, and (3) perform context reasoning with an MLLM to classify the new object as in- or out-of-context.
3.1 Inpainting
(a) Inpainting success rate.
(b) Context reasoning prompt.
Figure 3: (a) Inpainting success rate for original-inpainted object pairs. Rows are classes of original objects and columns are inpainted objects. A darker color indicates higher success rate. (b) Context reasoning prompt for Molmo [17].
We use Stable Diffusion 2 [47] for inpainting. Inpainting a new object requires an original image, an inpainting mask, and a prompt. For each COCO image, we randomly select an object. The object’s mask is slightly dilated and enclosed in a bounding box, which is used as the inpainting mask. In the preliminary study, we observed that the use of dilation and bounding boxes reduced residual artifacts from the original objects. For the replacement object, we randomly select one of COCO’s 80 object categories and use it as the text prompt.
The vanilla Stable Diffusion inpainting pipeline often struggles to inpaint small objects. To overcome this, we crop the original image around an enlarged inpainting area, inpaint, and then scale the results back to the original size. We use alpha blending to seamlessly merge the inpainted region with the original image. For more details on inpainting, see the supplementary materials.
3.2 Automated inpainting verification
Diffusion-based inpainting is imperfect: there is no guarantee the target object will be successfully inpainted. To confirm the success of inpainting, we use a COCO-trained YOLOv8x [48] object detector. If the inpainted object is detected within the inpainting mask, the process is deemed successful. For failed cases, we complete two additional rounds of inpainting and verification, pairing YOLOv8x with a state-of-the-art MLLM (Molmo [17]) to capture YOLO’s false negatives. Images that fail inpainting for all three rounds are discarded, resulting in 97,722 successful images.
Interestingly, we found that inpainting success rate is highly correlated with semantic priors. Figure 3(a) shows the inpainting success proportions for all original-inpainted object pairs. We note dense neighborhoods of inpainted object classes with increased probability of inpainting success. Using average linkage hierarchical clustering, we grouped replacement objects based on inpainting success proportions. With GloVe embeddings, we calculated cosine similarity between items within and across clusters. Mann-Whitney U tests revealed a significantly higher sample mean similarity within clusters (0.169) vs. across clusters (0.119) (bootstrapped p =.$=.002$). The cluster with the highest inpainting success included nearly all animals and food indicating a bias for these classes, while lowest performing cluster included furniture and household objects, which are highly context-dependent. A Kruskal-Wallis test confirmed significant differences across COCO supercategories (p $<0.001$), with food and animals outperforming electronics, appliances, and furniture. These findings validate the semantic coherence of the clusters formed and the vertical block pattern noted in Figure 3(a). Additional statistics are in the supplementary material. To our knowledge, this is the first thorough analysis of semantic priors in diffusion-based inpainting tasks.
3.3 Context reasoning
We label context for all inpainted objects. An object is considered in-context if it adheres to the contextual coherence of the scene with existing objects, while violations of the coherence result in an out-of-context classification. Our context reasoning is based on three fundamental principles regarding context  [2]: location, size, and co-occurrence. Location examines the spatial positioning of objects within a scene. Size evaluates whether the object’s dimensions are proportionate to the scene geometry. Co-occurrence measures whether objects’ simultaneous presence is common. To label context accordingly, we leverage the power of the Molmo [17]. We design a deliberate prompt to guide Molmo in performing language and visual reasoning for the context classification of our inpainted images (see Figure 3(b)). Provided with an inpainted image and context criteria, Molmo can conduct holistic visual-semantic reasoning that a pure language-based model cannot.
3.4 Comparison with prior datasets
Table 1: COinCO significantly enriches COCO with inpainting and context labels.
COCO (2017) COCO-Stuff COCOGlide COinCO
Dataset size 123K 164K 512 97,722
Unique source images 123K 164K 512 97,722
Replacement classes - - same as original 80
Out-of-context images - - -
Context reasoning - -
Inpainting - -
Labels objects objects, scenes (fake) objects (fake) objects, context
COinCO extends COCO by introducing contextual diversity with inpainted objects (Table 1). Unlike COCO and COCO-Stuff, which primarily focus on object and scene (stuff) annotations, COinCO provides intentional out-of-context scenarios, enabling research in contextual reasoning and fake localization. While COCOGlide also applies inpainting to COCO, it has significantly less images, and each replacement object has the same class as its original counterpart, thus the context is unaltered.
4 Manual Verification
To validate the reliability of our automated models for object detection and context classification, we conducted a comprehensive manual verification study. We manually annotated 1,000 images to validate the successful rate of our final dataset, and assessed whether successfully inpainted objects were in- or out-of-context, based on the aforementioned principles.
4.1 Verification on inpainted object detection
We first evaluated whether target objects can be reliably detected in the inpainted region. We compared COCO-trained detectors, including YOLO [48], MMDetection [49], and GroundingDINO [50] on 1,000 human-labeled images in Table 2. AR: Alignment Rate, PR: Precision, RE: Recall, FPR: False Positive Rate. AR is the primary metric reflecting the agreement between a model and a human.
Method AR (%) PR RE FPR
YOLOv8x 76.4 93.78 67.86 8.15
YOLOv11x 66.6 94.80 50.93 5.06
MMDet 67.6 86.87 58.54 16.01
GroundDINO 59.5 96.14 38.66 2.81
Table 2: Object detection performance.
Model Acc (%) Prec (%) Rec (%) F1 (%)
VisualNet 52.98 51.82 84.62 64.28
SemanticNet 76.33 78.19 73.02 75.52
VisSemanticNet 76.90 77.66 75.51 76.57
Table 3: Context classification performance.
For ground truth, annotators recorded whether the inpainted object was successfully integrated within the mask. We tuned object detector parameters, such as confidence, IoU threshold, and augmentation. YOLOv8x achieved the highest alignment rate (AR) in relation to the manual labels (76.4%)–AR is the rate at which the model agrees with human annotators. While GroundingDINO displayed marginally higher precision (96.14%) and a lower false positive rate (2.81%), the model’s recall was much lower than YOLOv8x (38.66% versus 67.86%), reflecting its failure to identify positives.
In addition, we tested Molmo [17] using cropped inpainting mask regions as input, prompting it to confirm whether the target object was present. Molmo achieved an AR of $82.6\%$ with human annotations, surpassing the performance of conventional object detectors. Thus, YOLOv8x and Molmo work together in our cascaded verification process for efficiency and accuracy.
4.2 Verification on context classification
To verify context classification, we established strict criteria. First, an agreement was required between Molmo and annotators regarding the presence of the target object (as in Section 4.1). Second, all human annotators were to agree on the context classification (in- or out-of-context). Of our 1000 manually annotated images, a total of 477 instances met these criteria. Among these instances, Molmo’s context classification aligned with human annotations in 370 cases ($77.57\%$). These results confirm Molmo’s reliability in context understanding and, together with YOLOv8x, validate our cascaded detection and classification verification process.
5 Analysis on COinCO
With COinCO as our testbed, we explore the role of context in several tasks. We demonstrate how our data can be used to learn a general model to classify objects as in- or out-of-context. We then propose a novel Objects-from-Context task for predicting which objects naturally belong in a given scene. Finally, we evaluate state-of-the-art fake detection models and show how contextual information can enhance fake localization without any fine-tuning. These applications demonstrate the versatility of COinCO and underscore the fundamental role of context in visual understanding. Examples of our pipeline are illustrated in Figure 4.
Original Mask Inpainted Fake Det. Context Reasoning Objects-from-Context
Refer to caption Refer to caption Refer to caption Refer to caption “A kite can reasonably be flying in the air above a railway track… The kite appears to be appropriately sized for flying, neither too large nor too small for the environment.
Final decision: In-context.
Refer to caption Refer to caption Refer to caption Refer to caption A cow should not be in the sky. It is impossible for a cow to appear together with the moon.
Final decision: Out-of-context.
Refer to caption Refer to caption Refer to caption Refer to caption “A horse in a bathroom, especially next to urinals, is highly unusual and unreasonable… Horses and urinals are completely unrelated objects that should not appear together…
Final decision: Out-of-context.
Refer to caption Refer to caption Refer to caption Refer to caption ”Cups and water are common together, but a cup floating on its own in a natural body of water is not reasonable.”
Final decision: Out-of-context.
Refer to caption Refer to caption Refer to caption Refer to caption “A soccer ball is reasonable in a bedroom setting…Soccer balls can appear together with beds, as they are both common items in a bedroom.
Final decision: In-context.
Refer to caption Refer to caption Refer to caption Refer to caption “A refrigerator is highly unusual and out of place in a zoo enclosure, especially one with giraffes and other animals…
Final decision: Out-of-context.”
Figure 4: Inpainting, fake detection, and objects-from-context results. Context reasoning responses are color-coded by location, size, and co-occurrence. Original objects are in red. Inpainted objects: kite, cow, horse, cup, sports ball, refrigerator, orange.
5.1 In- and out-of-context classification
COinCO’s context labels enables the training of a generalized context model to classify objects as in- or out-of-context within an image, addressing the question: “Does this object belong in this context?" We train a binary classification model that outputs a single score through a sigmoid activation function, where $1$: “out-of-context” and 00: “in-context.” For training and evaluation, we create a balanced dataset by supplementing our generated samples with original COCO images. As objects in COCO are naturally occurring, we assume they are “in-context”, ensuring a 50:50 ratio. We evaluate three models with different input modalities: (1) VisualNet: Uses only visual features from the image and mask via the VAE encoder in Stable Diffusion. (2) SemanticNet: Uses only semantic embeddings from BERT [51], including the average embedding of existing objects and the embedding of the query object. (3) VisSemanticNet: Uses both visual and semantic embeddings.
According to Table 3, the significant performance gap between VisualNet and SemanticNet shows that contextual understanding is primarily a reasoning task. VisualNet struggles to capture the complex relationships and real-world knowledge required for context reasoning and tends to predict objects as out-of-context, resulting in high recall but low accuracy and precision. VisSemanticNet achieves the best overall performance. Despite their simplicity, our context classification models show promising results and we hope they will inspire further research for more advanced context-aware models.
5.2 Objects-from-Context prediction
While our context classification model determines if an existing object belongs in a scene, our Objects-from-Context prediction task tackles a different question: “What object(s) fit to this context?” This task tests a model’s contextual comprehension of real images and aligns with the Context Challenge in [52]. We formulate this task at two levels: 1) Instance-level prediction. Predicting the exact object class among COCO’s 80 classes. 2) Clique-level prediction. Predicting the clique an object belongs to, using COCO’s super-categories: accessory, animal, appliance, electronic, food, furniture, indoor, kitchen, outdoor, person, sports, and vehicle.
We design a model that takes two inputs: the inpainted image and a binary mask indicating the target region. During training, we use the dilated bounding box of the original object as the mask and its class as the label. We use the VAE encoder from Stable Diffusion to extract latents from both inputs, which are processed by an MLP that predicts across COCO’s 80 classes. For clique-level evaluation, we map predicted object classes to their corresponding super-categories and check if they match the original objects. Importantly, our model can handle arbitrary mask sizes and locations during inference, making it adaptable for broader context reasoning applications. Detailed model architectures and training information are in the supplementary material. For comparison, we implement a baseline that ranks candidate objects based on the co-occurrence frequency of other objects in a scene [53]. If no other objects are present, this baseline defaults to random selection.
Table 4: Instance-level and clique-level object-from-context prediction accuracy.
Method Instance-level (%) Clique-level (%)
Top-1 Top-3 Top-5 Top-1 Top-3 Top-5
Random 1.25 3.75 6.25 8.33 25.00 41.67
Co-occurrence [53] 1.54 4.70 7.29 9.37 30.72 52.91
Ours 16.32 31.89 42.80 35.10 61.41 78.31
Table 4 shows our model’s superior grasp of contextual features, significantly outperforming random and co-occurrence baselines. At the instance level, the substantial improvement in top-3 and top-5 accuracy (31.89% and 42.80%) indicates that even when the model’s top prediction is incorrect, the true object is often ranked among the most probable candidates. Our model demonstrates stronger performance at the clique level, correctly predicting semantically similar objects. This result demonstrates that our approach effectively learns the relationship between scene contexts and the objects that belong within them. Our model offers three key applications as shown in the three rows of Figure 5: (1) analyzing suspicious regions to predict what objects existed before manipulation, serving as a reference for anomaly detection and forensic analysis [54]; (2) suggesting contextually appropriate objects based on spatial location within a scene, enabling intelligent image editing [33]; and (3) versatile analysis of any image (including unmodified images) to identify contextually coherent objects in regions of interest, supporting context-aware content generation. This versatility advances context-driven visual understanding across multiple practical applications.
Refer to caption Refer to caption Refer to caption
Instance P(%) skis: 31.6 snowboard: 24.3 kite: 14.8 surfboard: 7.0 backpack: 5.9 Clique P(%) sports: 79.2 accessory: 6.6 person: 4.9 vehicle: 4.2 animal: 3.0 Instance P(%) kite: 22.9 skis: 22.1 snowboard: 19.6 surfboard: 8.6 person: 5.5 Clique P(%) sports: 75.1 accessory: 6.1 vehicle: 5.7 person: 5.5 animal: 4.2 Instance P(%) horse: 5.0 zebra: 5.0 giraffe: 4.9 person: 4.6 elephant: 4.4 Clique P(%) animal: 37.2 vehicle: 14.9 sports: 12.1 outdoor: 10.7 accessory: 6.8
Figure 5: Object-from-Context prediction. A red box is a query. The top row shows three examples (two inpainted, one original COCO), and the bottom row lists instance-level and clique-level predictions with their probabilities (P). Objects in red are the top predictions.
Method Acc F1 AUC AP
ManTraNet [42] 85.4 30.7 84.9 50.7
Trufor [43] 89.7 55.9 93.4 73.6
PSCC-Net [40] 89.5 46.8 95.4 79.5
CAT-Net [41] 92.7 76.5 97.4 90.3
Table 5: SOTA fake detection performance.
Figure 6: Context enhancement results.
5.3 Context-empowered fake localization
Fake localization aims to identify the specific synthetic regions in an image, producing a pixel-level map that is fine-grained compared to binary image-level classification. In COinCO, the ground truth mask for fake localization is the original object’s entire bounding box used in inpainting. We evaluate localization performance using four common pixel-level metrics [41, 40]: F1 Score and Accuracy (threshold of $0.5$), Area Under Curve (AUC), and Average Precision (AP).
State-of-the-art performance. We benchmark several SOTA fake detection models on COinCO (Table 5). The high accuracy and AUC scores ($>84\%$ for all models) can be attributed to the small average size of inpainted regions in our data. As most images are majority authentic, these metrics are biased towards high values. F1 and AP, which better reflect the precision of fake localization, reveal a clear performance gap among models. CAT-Net achieves the best performance across all metrics with $92.65\%$ accuracy, $76.47\%$ F1 score, $97.35\%$ AUC, and $90.33\%$ AP. Trufor obtains the second-best F1 score of $55.92\%$, while PSCC-Net shows strong performance in AUC ($95.38\%$) and AP ($79.52\%$). ManTraNet demonstrates relatively lower performance, particularly in F1 score ($30.68\%$) and AP ($50.72\%$). These findings highlight significant differences and room for improvement among SOTA fake localizers. Examples of these SOTA models’ predictions are shown in Figure 7.
Incorporating context in fake localization. To leverage contextual information for improved fake localization, we propose a simple yet effective way of enhancing the prediction scores of fake localization models in regions corresponding to out-of-context objects. Concretely, the context-enhanced prediction score $P^{\prime}(x,y)$ at pixel $(x,y)$ is defined as $\min(P(x,y)\times\gamma,\,1.0),\text{if }(x,y)\in M_{\text{OC}}$; otherwise $\text{otherwise as}P(x,y)$, the original predicted fake score, $M_{\text{OC}}$ is the mask region of out-of-context object (predicted by context reasoning), $\gamma$ is a context enhancement factor, and the $\min$ function ensures the enhanced score $\leq 1.0$.
Image GT [41] [42] [43] [40]
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 7: Fake localization performance.
We evaluate this context-enhancement approach under two settings. In the oracle setting, we use ground truth annotations to enhance predictions within the fake object’s mask region when it is out-of-context. This serves as an upper bound for performance gains. In the practical setting, the fake objects are unknown. We propose the use of Molmo to identify suspicious objects based on size and location, deliberately excluding co-occurrence to prevent false positives. For instance, if an image contains only two objects, e.g., an apple and an inpainted traffic light, either might be seen as out-of-context. By focusing on physical violations, our method remains robust while effectively detecting fake objects. As Molmo may identify multiple out-of-context objects in an image, we define $M_{\text{OC}}$ as the union of these objects’ masks. In both settings, we set the enhancement factor $\gamma$ to $5$ to enhance prediction scores in out-of-context regions. A detailed analysis of $\gamma$ is in supplementary materials.
Figure 6 shows how our context-enhancement improves fake localization for all SOTA detectors without any finetuning. The effectiveness of our method stems from the complementary nature of context detection and fake localization. This combination ensures that even if Molmo misidentifies authentic objects as out-of-context, enhancement only strengthens the base model’s predictions, preserving robustness. These results highlight context as a valuable signal for fake detection, opening new possibilities for context-aware image forensics.
6 Conclusion and Limitations
We present COinCO, a novel dataset designed to feature in- and out-of-context objects in real scenes. By strategically replacing objects in COCO with diffusion-based inpainting, we systematically diversify the contextual status of objects in complex scenes. COinCO’s context labels allow us to train a general context classifier. We also introduce a novel Object-from-Context prediction task. Finally, our work advances current fake detection approaches by leveraging contextual information during fake pixel localization. COinCO and our findings highlight the importance of context in visual understanding and provide a foundation and testbed for future research.
While COinCO advances context-aware visual understanding, context classification remains inherently subjective despite our structured criteria, as human annotators and MLLMs occasionally disagree on contextual coherence. The Objects-from-Context prediction task is currently limited to COCO’s predefined categories. Future work could explore extending these tasks to open-vocabulary settings for more flexible context reasoning.